OpenStudy (anonymous):

The perimeter of a regular hexagon is seventy-two and each side is represented by three x. What is x? 4 6 12 36

OpenStudy (anonymous):

@amistre64 @Compassionate @ganeshie8 @TexasBoy123 @undeadknight26

OpenStudy (anonymous):

4??

jhonyy9 (jhonyy9):

hexagon mean six sides then i understand it right 3x is a side ,yes ?

OpenStudy (anonymous):

ok thank you!

jhonyy9 (jhonyy9):

so then 3x*6 =72 18x=72 x=72/18 =4 x=4

jhonyy9 (jhonyy9):

yes are right